Solve by applying the Quadratic Formula; all radicals should be simplified as far as possible. Show your work.
12d^2 + 29d + 15 = 0

Answer by checkley77(12844)

12d^2 + 29d + 15 = 0
(4d+3)(3d+5)=0
4d+3=0
4d=-3
d=-3/4 ans.
3d+5=0
3d=-5
d=-5/3 ans.